OpenStudy (anonymous):

WILL GIVE MEDAL! What is the solution set of y = x2 + 2x + 7 and y = x + 7? {(0, 7), (-1, 6)} {(0, 7), (-7, 0)} {(0, 7), (1, 8)} {(-2, 0), (4, 0)} ***Its NOT B***

OpenStudy (anonymous):

have you tried solving it? \(\sf y = x^2 + 2x + 7\) and \(\sf \color{red}{y = x + 7}\) plug the value of y in red into the quadratic equation and solve for x \(\sf \color{red}{x + 7} = x^2 + 2x + 7\)

OpenStudy (anonymous):

Yea I did. I dont understand it at all.

OpenStudy (anonymous):

can you show me your solution?

OpenStudy (anonymous):

I dont have it anymore it was in a different question

OpenStudy (anonymous):

well then, can you continue what I started above?

OpenStudy (anonymous):

No Im so confused

OpenStudy (anonymous):

which part are you confused with?

OpenStudy (anonymous):

How to solve

OpenStudy (anonymous):

@wiggles12

OpenStudy (anonymous):

read the first post I posted, that's just the first step we're combining the two functions together to find their solutions: